# Electric Dipole Moments in 5+3 Flavor Weak Effective Theory

### v1

@@ 5.2 Master formula for $\omega_{\text{HfF}}$
$[C_{e\gamma}]_{11}$ $(-28.7\pm 1.1)\cdot 10^{11}$ $[C^{S,RR}_{eu}]_{1122}$ $(93.0\pm 3.1)\cdot 10^{4}$
$[C^{S,RR}_{ee}]_{1111}$ $(-27.0\pm 1.1)\cdot 10^{3}$ $[C^{T,RR}_{eu}]_{1122}$ $(-38.4\pm 1.4)\cdot 10^{7}$
$[C^{S,RR}_{eu}]_{1111}$ $(-25.7\pm 2.4)\cdot 10^{7}$ $[C^{S,RR}_{ed}]_{1133}$ $(33.2\pm 1.3)\cdot 10^{4}$
$[C^{S,RR}_{ed}]_{1111}$ $(-26.1\pm 2.8)\cdot 10^{7}$ $[C^{T,RR}_{ed}]_{1133}$ $(46.4\pm 2.2)\cdot 10^{7}$
$[C^{S,RR}_{ed}]_{1122}$ $(-19.4\pm 1.5)\cdot 10^{6}$ $[C^{S,RR}_{ee}]_{1122}$ $(58.4\pm 2.3)\cdot 10^{3}$
$[C^{S,RL}_{eu}]_{1111}$ $(-25.6\pm 2.7)\cdot 10^{7}$ $[C^{S,RR}_{ee}]_{1133}$ $(127.2\pm 5.7)\cdot 10^{4}$
$[C^{S,RL}_{ed}]_{1111}$ $(-26.1\pm 2.2)\cdot 10^{7}$ $[C^{S,RR}_{ee}]_{1221}$ $(-28.5\pm 1.0)\cdot 10^{5}$
$[C^{S,RL}_{ed}]_{1122}$ $(-19.2\pm 1.5)\cdot 10^{6}$ $[C^{S,RR}_{ee}]_{1331}$ $(-55.2\pm 2.4)\cdot 10^{6}$
$[C^{T,RR}_{eu}]_{1111}$ $(232.4\pm 8.6)\cdot 10^{6}$ $[C^{V,LR}_{ee}]_{1221}$ $(72.7\pm 3.2)\cdot 10^{3}$
$[C^{T,RR}_{ed}]_{1111}$ $(-116.3\pm 5.1)\cdot 10^{6}$ $[C^{V,LR}_{ee}]_{1331}$ $(72.7\pm 3.0)\cdot 10^{3}$
$[C^{T,RR}_{ed}]_{1122}$ $(-89.9\pm 3.6)\cdot 10^{6}$
Table 7: Master formula for $\omega_{\rm HfF}$ in WET in the JMS basis at the EW scale. The $\alpha_{I}^{\rm HfF}$ have units of (mrad/s)$\rm TeV^{2}$ and (mrad/s)$\rm TeV$ for the 4-Fermi and dipole operators, respectively. The $90\%$ CL upper bound on $\omega_{\rm HfF}$ is $0.17$ mrad/s.
@@ 5.3 Master formula for $d_{\rm Hg}$
$\alpha_{I}^{\rm Hg}(\mu_{ew})$ for the $d_{\text{Hg}}$: SL operators
$n_{q}+n_{\ell}=5+3$
$[C^{S,RR}_{eu}]_{1111}$ $(8.3\pm 1.5)\cdot 10^{-22}$
$[C^{S,RR}_{ed}]_{1111}$ $(-3.8\pm 1.4)\cdot 10^{-22}$
$[C^{S,RR}_{ed}]_{1122}$ $(16.8\pm 3.7)\cdot 10^{-24}$
$[C^{S,RL}_{eu}]_{1111}$ $(-3.8\pm 1.3)\cdot 10^{-22}$
$[C^{S,RL}_{ed}]_{1111}$ $(8.3\pm 1.5)\cdot 10^{-22}$
$[C^{S,RL}_{ed}]_{1122}$ $(16.8\pm 3.8)\cdot 10^{-24}$
$[C^{T,RR}_{eu}]_{1111}$ $(-3.7\pm 1.2)\cdot 10^{-23}$
$[C^{T,RR}_{ed}]_{1111}$ $(-24.1\pm 5.8)\cdot 10^{-23}$
$[C^{T,RR}_{ed}]_{1122}$ $(17.3\pm 4.7)\cdot 10^{-25}$
$[C^{T,RR}_{eu}]_{1122}$ $(-3.0\pm 1.8)\cdot 10^{-33}$
$[C^{S,RR}_{eu}]_{1122}$
$[C^{T,RR}_{ed}]_{1133}$
$[C^{T,RR}_{ed}]_{2211}$ $(-25.7\pm 6.4)\cdot 10^{-26}$
$[C^{T,RR}_{ed}]_{2222}$ $(10.0\pm 7.1)\cdot 10^{-28}$
$[C^{T,RR}_{ed}]_{2233}$ $(12.6\pm 6.9)\cdot 10^{-33}$
$[C^{T,RR}_{ed}]_{3311}$ $(-4.7\pm 1.1)\cdot 10^{-24}$
$[C^{T,RR}_{ed}]_{3322}$ $(18.4\pm 9.6)\cdot 10^{-27}$
$[C^{T,RR}_{ed}]_{3333}$ $(2.1\pm 1.5)\cdot 10^{-31}$
$[C^{T,RR}_{eu}]_{2211}$ $(4.1\pm 1.4)\cdot 10^{-26}$
$[C^{T,RR}_{eu}]_{2222}$ $(-6.2\pm 3.6)\cdot 10^{-31}$
$[C^{T,RR}_{eu}]_{3311}$ $(7.6\pm 2.6)\cdot 10^{-25}$
$[C^{T,RR}_{eu}]_{3322}$ $(-10.4\pm 5.9)\cdot 10^{-30}$
$[C^{S,RR}_{ed}]_{1133}$
$[C^{S,RR}_{ed}]_{2211}$ $(-25.4\pm 6.4)\cdot 10^{-29}$
$[C^{S,RR}_{ed}]_{2222}$ $(9.9\pm 5.7)\cdot 10^{-31}$
$[C^{S,RR}_{ed}]_{2233}$
$[C^{S,RR}_{ed}]_{3311}$ $(-5.2\pm 1.4)\cdot 10^{-27}$
$[C^{S,RR}_{ed}]_{3322}$ $(2.0\pm 1.2)\cdot 10^{-29}$
$[C^{S,RR}_{ed}]_{3333}$
$[C^{S,RR}_{eu}]_{2211}$ $(-8.1\pm 3.0)\cdot 10^{-29}$
$[C^{S,RR}_{eu}]_{2222}$
$[C^{S,RR}_{eu}]_{3311}$ $(-16.5\pm 6.6)\cdot 10^{-28}$
$[C^{S,RR}_{eu}]_{3322}$ $(7.1\pm 3.7)\cdot 10^{-33}$
@@
$C_{\widetilde{G}}$ $(-1.9\pm 1.1)\cdot 10^{-25}$ $[C_{uG}]_{22}$ $(-3.7\pm 1.9)\cdot 10^{-24}$ $[C_{dG}]_{33}$ $(-2.6\pm 1.4)\cdot 10^{-25}$
$[C_{dG}]_{11}$ $(13.7\pm 8.0)\cdot 10^{-21}$ $[C_{u\gamma}]_{22}$ $(1.8\pm 1.0)\cdot 10^{-25}$ $[C_{d\gamma}]_{33}$ $(-4.8\pm 2.7)\cdot 10^{-27}$
$[C_{dG}]_{22}$ $(-6.7\pm 4.1)\cdot 10^{-24}$ $[C^{S1,RR}_{uu}]_{2222}$ $(2.7\pm 1.6)\cdot 10^{-28}$ $[C^{V1,LR}_{uddu}]_{1331}$ $(-1.9\pm 1.5)\cdot 10^{-26}$
$[C_{uG}]_{11}$ $(5.3\pm 2.6)\cdot 10^{-21}$ $[C^{S8,RR}_{uu}]_{2222}$ $(-6.1\pm 3.2)\cdot 10^{-29}$ $[C^{V8,LR}_{uddu}]_{1331}$ $(-1.4\pm 1.0)\cdot 10^{-26}$
$[C_{d\gamma}]_{11}$ $(34.5\pm 9.5)\cdot 10^{-21}$ $[C^{S1,RR}_{uu}]_{1122}$ $(6.9\pm 3.9)\cdot 10^{-26}$ $[C^{V1,LR}_{dd}]_{1331}$ $(7.8\pm 2.8)\cdot 10^{-26}$
$[C_{d\gamma}]_{22}$ $(-13.4\pm 9.0)\cdot 10^{-23}$ $[C^{S8,RR}_{uu}]_{1122}$ $(1.3\pm 1.0)\cdot 10^{-26}$ $[C^{V8,LR}_{dd}]_{1331}$ $(20.0\pm 5.0)\cdot 10^{-26}$
$[C_{u\gamma}]_{11}$ $(-5.6\pm 1.8)\cdot 10^{-21}$ $[C^{S1,RR}_{ud}]_{2211}$ $(1.8\pm 1.0)\cdot 10^{-25}$ $[C^{V1,LR}_{dd}]_{2332}$ $(-4.7\pm 3.0)\cdot 10^{-28}$
$[C^{V1,LR}_{uddu}]_{1111}$ $(1.3\pm 5.2)\cdot 10^{-23}$ $[C^{S8,RR}_{ud}]_{2211}$ $(10.7\pm 3.9)\cdot 10^{-26}$ $[C^{V8,LR}_{dd}]_{2332}$ $(-6.3\pm 4.5)\cdot 10^{-28}$
$[C^{V8,LR}_{uddu}]_{1111}$ $(2.1\pm 7.9)\cdot 10^{-23}$ $[C^{S1,RR}_{ud}]_{2222}$ $(-12.6\pm 5.5)\cdot 10^{-30}$ $[C^{S1,RR}_{dd}]_{3333}$ $(5.0\pm 2.7)\cdot 10^{-29}$
$[C^{V1,LR}_{uddu}]_{1221}$ $(0.7\pm 2.5)\cdot 10^{-23}$ $[C^{S8,RR}_{ud}]_{2222}$ $(-2.3\pm 1.5)\cdot 10^{-28}$ $[C^{S8,RR}_{dd}]_{3333}$ $(-10.3\pm 5.0)\cdot 10^{-30}$
$[C^{V8,LR}_{uddu}]_{1221}$ $(1.1\pm 4.2)\cdot 10^{-23}$ $[C^{S1,RR}_{uddu}]_{2112}$ $(-9.8\pm 4.4)\cdot 10^{-25}$ $[C^{S1,RR}_{dd}]_{1133}$ $(2.4\pm 1.3)\cdot 10^{-25}$
$[C^{V1,LR}_{dd}]_{1221}$ $(-0.6\pm 2.3)\cdot 10^{-23}$ $[C^{S8,RR}_{uddu}]_{2112}$ $(-15.1\pm 4.7)\cdot 10^{-26}$ $[C^{S8,RR}_{dd}]_{1133}$ $(3.4\pm 4.1)\cdot 10^{-26}$
$[C^{V8,LR}_{dd}]_{1221}$ $(-1.0\pm 3.8)\cdot 10^{-23}$ $[C^{S1,RR}_{uddu}]_{2222}$ $(11.1\pm 7.1)\cdot 10^{-28}$ $[C^{S1,RR}_{dd}]_{2233}$ $(-5.5\pm 3.2)\cdot 10^{-29}$
$[C^{S1,RR}_{uu}]_{1111}$ $(-0.6\pm 1.7)\cdot 10^{-23}$ $[C^{S8,RR}_{uddu}]_{2222}$ $(8.5\pm 5.4)\cdot 10^{-28}$ $[C^{S8,RR}_{dd}]_{2233}$ $(1.6\pm 1.2)\cdot 10^{-28}$
$[C^{S8,RR}_{uu}]_{1111}$ $(1.8\pm 6.2)\cdot 10^{-24}$ $[C^{S1,RR}_{uu}]_{1221}$ $(-2.5\pm 1.5)\cdot 10^{-25}$ $[C^{S1,RR}_{dd}]_{1331}$ $(-13.0\pm 8.1)\cdot 10^{-25}$
$[C^{S1,RR}_{dd}]_{1111}$ $(0.5\pm 1.6)\cdot 10^{-23}$ $[C^{S8,RR}_{uu}]_{1221}$ $(6.5\pm 2.1)\cdot 10^{-26}$ $[C^{S8,RR}_{dd}]_{1331}$ $(5.1\pm 1.7)\cdot 10^{-25}$
$[C^{S8,RR}_{dd}]_{1111}$ $(-1.5\pm 6.0)\cdot 10^{-24}$ $[C^{V1,LR}_{uddu}]_{2112}$ $(-34.4\pm 9.0)\cdot 10^{-27}$ $[C^{S1,RR}_{dd}]_{2332}$ $(-6.5\pm 4.6)\cdot 10^{-28}$
$[C^{S1,RR}_{dd}]_{2222}$ $(-4.5\pm 2.9)\cdot 10^{-29}$ $[C^{V8,LR}_{uddu}]_{2112}$ $(-8.3\pm 2.2)\cdot 10^{-26}$ $[C^{S8,RR}_{dd}]_{2332}$ $(-12.7\pm 9.3)\cdot 10^{-28}$
$[C^{S8,RR}_{dd}]_{2222}$ $(-5.3\pm 3.5)\cdot 10^{-29}$ $[C^{V1,LR}_{uddu}]_{2222}$ $(9.4\pm 5.5)\cdot 10^{-29}$ $[C^{S1,RR}_{ud}]_{1133}$ $(9.3\pm 5.7)\cdot 10^{-26}$
$[C^{S1,RR}_{dd}]_{1122}$ $(1.2\pm 5.5)\cdot 10^{-24}$ $[C^{V8,LR}_{uddu}]_{2222}$ $(3.7\pm 2.8)\cdot 10^{-28}$ $[C^{S8,RR}_{ud}]_{1133}$ $(3.8\pm 1.8)\cdot 10^{-26}$
$[C^{S8,RR}_{dd}]_{1122}$ $(-0.4\pm 1.4)\cdot 10^{-24}$ $[C^{V1,LR}_{uu}]_{1221}$ $(5.4\pm 3.1)\cdot 10^{-27}$ $[C^{S1,RR}_{uddu}]_{1331}$ $(-6.7\pm 3.5)\cdot 10^{-25}$
$[C^{S1,RR}_{dd}]_{1221}$ $(1.2\pm 5.8)\cdot 10^{-24}$ $[C^{V8,LR}_{uu}]_{1221}$ $(18.5\pm 6.4)\cdot 10^{-27}$ $[C^{S8,RR}_{uddu}]_{1331}$ $(2.8\pm 3.9)\cdot 10^{-26}$
$[C^{S8,RR}_{dd}]_{1221}$ $(-0.4\pm 1.7)\cdot 10^{-24}$ $[C^{S1,RR}_{ud}]_{2233}$ $(-14.2\pm 8.1)\cdot 10^{-29}$
$[C^{S1,RR}_{ud}]_{1111}$ $(-2.3\pm 4.1)\cdot 10^{-25}$ $[C^{S8,RR}_{ud}]_{2233}$ $(-4.3\pm 2.9)\cdot 10^{-29}$
$[C^{S8,RR}_{ud}]_{1111}$ $(0.7\pm 1.2)\cdot 10^{-25}$ $[C^{S1,RR}_{uddu}]_{2332}$ $(6.8\pm 4.0)\cdot 10^{-28}$
$[C^{S1,RR}_{ud}]_{1122}$ $(-1.5\pm 4.0)\cdot 10^{-24}$ $[C^{S8,RR}_{uddu}]_{2332}$ $(-7.3\pm 4.4)\cdot 10^{-29}$
$[C^{S8,RR}_{ud}]_{1122}$ $(0.4\pm 1.3)\cdot 10^{-24}$ $[C^{V1,LR}_{uddu}]_{2332}$ $(18.2\pm 9.9)\cdot 10^{-30}$
$[C^{S1,RR}_{uddu}]_{1111}$ $(-2.3\pm 3.8)\cdot 10^{-25}$ $[C^{V8,LR}_{uddu}]_{2332}$ $(-12.0\pm 6.8)\cdot 10^{-30}$
$[C^{S8,RR}_{uddu}]_{1111}$ $(6.8\pm 9.3)\cdot 10^{-26}$
$[C^{S1,RR}_{uddu}]_{1221}$ $(-1.5\pm 4.5)\cdot 10^{-24}$
$[C^{S8,RR}_{uddu}]_{1221}$ $(0.4\pm 1.3)\cdot 10^{-24}$
@@
Table 9: Master Formula for $d_{\rm Hg}$ in WET in the JMS basis at the EW scale. The $\alpha_{I}^{\rm Hg}$ have units of $e$ cm $\rm TeV^{2}$ and $e$ cm $\rm TeV$ for the 4-Fermi and dipole operators, respectively. The $90\%$ CL upper bound on $d_{\rm Hg}$ is $6.2\times 10^{-30}$ $e$ cm.

### v2

@@ 5.2 $\omega_{\text{HfF}}$
$[C_{e\gamma}]_{11}$ $(-2.9\pm 0.1)\cdot 10^{12}$ $[C^{S,RR}_{eu}]_{1122}$ $(9.2\pm 0.3)\cdot 10^{5}$
$[C^{S,RR}_{ee}]_{1111}$ $(-2.7\pm 0.1)\cdot 10^{4}$ $[C^{T,RR}_{eu}]_{1122}$ $(-3.8\pm 0.2)\cdot 10^{8}$
$[C^{S,RR}_{eu}]_{1111}$ $(-2.6\pm 0.2)\cdot 10^{8}$ $[C^{S,RR}_{ed}]_{1133}$ $(3.7\pm 0.1)\cdot 10^{5}$
$[C^{S,RR}_{ed}]_{1111}$ $(-2.6\pm 0.3)\cdot 10^{8}$ $[C^{T,RR}_{ed}]_{1133}$ $(4.8\pm 0.2)\cdot 10^{8}$
$[C^{S,RR}_{ed}]_{1122}$ $(-1.9\pm 0.1)\cdot 10^{7}$ $[C^{S,RR}_{ee}]_{1122}$ $(5.8\pm 0.2)\cdot 10^{4}$
$[C^{S,RL}_{eu}]_{1111}$ $(-2.6\pm 0.3)\cdot 10^{8}$ $[C^{S,RR}_{ee}]_{1133}$ $(1.3\pm 0.0)\cdot 10^{6}$
$[C^{S,RL}_{ed}]_{1111}$ $(-2.6\pm 0.2)\cdot 10^{8}$ $[C^{S,RR}_{ee}]_{1221}$ $(-2.8\pm 0.1)\cdot 10^{6}$
$[C^{S,RL}_{ed}]_{1122}$ $(-1.9\pm 0.1)\cdot 10^{7}$ $[C^{S,RR}_{ee}]_{1331}$ $(-5.5\pm 0.2)\cdot 10^{7}$
$[C^{T,RR}_{eu}]_{1111}$ $(2.3\pm 0.1)\cdot 10^{8}$ $[C^{V,LR}_{ee}]_{1221}$ $(9.1\pm 0.4)\cdot 10^{5}$
$[C^{T,RR}_{ed}]_{1111}$ $(-1.2\pm 0.0)\cdot 10^{8}$ $[C^{V,LR}_{ee}]_{1331}$ $(1.5\pm 0.1)\cdot 10^{7}$
$[C^{T,RR}_{ed}]_{1122}$ $(-9.0\pm 0.4)\cdot 10^{7}$
Table 7: Master formula for $\omega_{\rm HfF}$ in WET in the JMS basis at the EW scale. The $\alpha_{I}^{\rm HfF}$ coefficients have units of (mrad/s)$\rm TeV^{2}$ and (mrad/s)$\rm TeV$ for the four-fermion and dipole operators, respectively. The $90\%$ CL upper bound on $\omega_{\rm HfF}$ is $0.17$ mrad/s.
@@ 5.3 $d_{\rm Hg}$
$\alpha_{I}^{\rm Hg}(\mu_{ew})$ for the Hg EDM: SL operators
$n_{f}=3+1$ $n_{f}=4+1$ $n_{f}=5+3$
$[C^{S,RR}_{eu}]_{1111}$ $(8.3\pm 1.5)\cdot 10^{-22}$ $[C^{T,RR}_{eu}]_{1122}$ $[C^{T,RR}_{ed}]_{1133}$
$[C^{S,RR}_{ed}]_{1111}$ $(-3.8\pm 1.3)\cdot 10^{-22}$ $[C^{S,RR}_{eu}]_{1122}$ $[C^{T,RR}_{ed}]_{2211}$ $(-2.6\pm 0.7)\cdot 10^{-25}$
$[C^{S,RR}_{ed}]_{1122}$ $(1.7\pm 0.4)\cdot 10^{-23}$ $[C^{T,RR}_{ed}]_{2222}$ $(1.0\pm 0.7)\cdot 10^{-27}$
$[C^{S,RL}_{eu}]_{1111}$ $(-3.8\pm 1.4)\cdot 10^{-22}$ $[C^{T,RR}_{ed}]_{2233}$
$[C^{S,RL}_{ed}]_{1111}$ $(8.3\pm 1.6)\cdot 10^{-22}$ $[C^{T,RR}_{ed}]_{3311}$ $(-4.8\pm 1.4)\cdot 10^{-24}$
$[C^{S,RL}_{ed}]_{1122}$ $(1.7\pm 0.4)\cdot 10^{-23}$ $[C^{T,RR}_{ed}]_{3322}$ $(1.9\pm 1.3)\cdot 10^{-26}$
$[C^{T,RR}_{eu}]_{1111}$ $(-3.7\pm 0.9)\cdot 10^{-23}$ $[C^{T,RR}_{ed}]_{3333}$ $(3.5\pm 2.1)\cdot 10^{-31}$
$[C^{T,RR}_{ed}]_{1111}$ $(-2.4\pm 0.6)\cdot 10^{-22}$ $[C^{T,RR}_{eu}]_{2211}$ $(5.0\pm 8.0)\cdot 10^{-26}$
$[C^{T,RR}_{ed}]_{1122}$ $(1.7\pm 0.6)\cdot 10^{-24}$ $[C^{T,RR}_{eu}]_{2222}$ $(-5.9\pm 3.8)\cdot 10^{-31}$
- - $[C^{T,RR}_{eu}]_{3311}$ $(9.1\pm 12.1)\cdot 10^{-25}$
- - $[C^{T,RR}_{eu}]_{3322}$ $(-9.9\pm 5.7)\cdot 10^{-30}$
- - $[C^{S,RR}_{ed}]_{1133}$
- - $[C^{S,RR}_{ed}]_{2211}$ $(-2.6\pm 0.6)\cdot 10^{-28}$
- - $[C^{S,RR}_{ed}]_{2222}$ $(1.0\pm 0.6)\cdot 10^{-30}$
- - $[C^{S,RR}_{ed}]_{2233}$
- - $[C^{S,RR}_{ed}]_{3311}$ $(-5.2\pm 1.4)\cdot 10^{-27}$
- - $[C^{S,RR}_{ed}]_{3322}$ $(2.0\pm 1.3)\cdot 10^{-29}$
- - $[C^{S,RR}_{ed}]_{3333}$
- - $[C^{S,RR}_{eu}]_{2211}$ $(-8.9\pm 5.2)\cdot 10^{-29}$
- - $[C^{S,RR}_{eu}]_{2222}$
- - $[C^{S,RR}_{eu}]_{3311}$ $(-1.8\pm 1.0)\cdot 10^{-27}$
- - $[C^{S,RR}_{eu}]_{3322}$
@@
$C_{\widetilde{G}}$ $(-2.4\pm 3.9)\cdot 10^{-25}$ $[C_{uG}]_{22}$ $(-2.6\pm 1.5)\cdot 10^{-24}$ $[C_{dG}]_{33}$ $(-3.8\pm 2.6)\cdot 10^{-25}$
$[C_{dG}]_{11}$ $(5.0\pm 28.3)\cdot 10^{-20}$ $[C_{u\gamma}]_{22}$ $(2.5\pm 1.6)\cdot 10^{-25}$ $[C_{d\gamma}]_{33}$ $(-7.3\pm 5.3)\cdot 10^{-27}$
$[C_{dG}]_{22}$ $(-6.7\pm 4.9)\cdot 10^{-24}$ $[C^{S1,RR}_{uu}]_{2222}$ $(1.9\pm 1.0)\cdot 10^{-28}$ $[C^{V1,LR}_{uddu}]_{1331}$ $(-1.4\pm 7.4)\cdot 10^{-24}$
$[C_{uG}]_{11}$ $(4.1\pm 32.9)\cdot 10^{-20}$ $[C^{S8,RR}_{uu}]_{2222}$ $(-4.3\pm 2.6)\cdot 10^{-29}$ $[C^{V8,LR}_{uddu}]_{1331}$ $(3.2\pm 21.5)\cdot 10^{-25}$
$[C_{d\gamma}]_{11}$ $(3.5\pm 1.3)\cdot 10^{-20}$ $[C^{S1,RR}_{uu}]_{1122}$ $(5.9\pm 46.4)\cdot 10^{-25}$ $[C^{V1,LR}_{dd}]_{1331}$ $(-1.6\pm 7.4)\cdot 10^{-24}$
$[C_{d\gamma}]_{22}$ $(-1.3\pm 0.8)\cdot 10^{-22}$ $[C^{S8,RR}_{uu}]_{1122}$ $(1.7\pm 15.9)\cdot 10^{-25}$ $[C^{V8,LR}_{dd}]_{1331}$ $(9.5\pm 20.4)\cdot 10^{-25}$
$[C_{u\gamma}]_{11}$ $(-9.4\pm 21.8)\cdot 10^{-21}$ $[C^{S1,RR}_{ud}]_{2211}$ $(7.0\pm 35.1)\cdot 10^{-25}$ $[C^{V1,LR}_{dd}]_{2332}$ $(-5.0\pm 3.3)\cdot 10^{-28}$
$[C^{V1,LR}_{uddu}]_{1111}$ $(8.9\pm 30.3)\cdot 10^{-24}$ $[C^{S8,RR}_{ud}]_{2211}$ $(2.6\pm 12.5)\cdot 10^{-25}$ $[C^{V8,LR}_{dd}]_{2332}$ $(-1.8\pm 1.1)\cdot 10^{-27}$
$[C^{V8,LR}_{uddu}]_{1111}$ $(1.4\pm 5.3)\cdot 10^{-23}$ $[C^{S1,RR}_{ud}]_{2222}$ $(-1.2\pm 0.6)\cdot 10^{-29}$ $[C^{S1,RR}_{dd}]_{3333}$ $(7.8\pm 5.0)\cdot 10^{-29}$
$[C^{V1,LR}_{uddu}]_{1221}$ $(4.8\pm 14.4)\cdot 10^{-24}$ $[C^{S8,RR}_{ud}]_{2222}$ $(-2.3\pm 1.4)\cdot 10^{-28}$ $[C^{S8,RR}_{dd}]_{3333}$ $(-1.6\pm 1.2)\cdot 10^{-29}$
$[C^{V8,LR}_{uddu}]_{1221}$ $(7.4\pm 23.5)\cdot 10^{-24}$ $[C^{S1,RR}_{uddu}]_{2112}$ $(-3.1\pm 15.0)\cdot 10^{-24}$ $[C^{S1,RR}_{dd}]_{1133}$ $(9.1\pm 52.2)\cdot 10^{-25}$
$[C^{V1,LR}_{dd}]_{1221}$ $(-4.1\pm 13.9)\cdot 10^{-24}$ $[C^{S8,RR}_{uddu}]_{2112}$ $(5.1\pm 201.5)\cdot 10^{-26}$ $[C^{S8,RR}_{dd}]_{1133}$ $(2.6\pm 14.4)\cdot 10^{-25}$
$[C^{V8,LR}_{dd}]_{1221}$ $(-6.2\pm 20.0)\cdot 10^{-24}$ $[C^{S1,RR}_{uddu}]_{2222}$ $(1.1\pm 0.6)\cdot 10^{-27}$ $[C^{S1,RR}_{dd}]_{2233}$ $(-5.6\pm 3.5)\cdot 10^{-29}$
$[C^{S1,RR}_{uu}]_{1111}$ $(-5.9\pm 21.7)\cdot 10^{-24}$ $[C^{S8,RR}_{uddu}]_{2222}$ $(8.5\pm 6.3)\cdot 10^{-28}$ $[C^{S8,RR}_{dd}]_{2233}$ $(1.6\pm 1.1)\cdot 10^{-28}$
$[C^{S8,RR}_{uu}]_{1111}$ $(1.8\pm 5.6)\cdot 10^{-24}$ $[C^{S1,RR}_{uu}]_{1221}$ $(-2.4\pm 15.3)\cdot 10^{-24}$ $[C^{S1,RR}_{dd}]_{1331}$ $(-5.7\pm 34.0)\cdot 10^{-24}$
$[C^{S1,RR}_{dd}]_{1111}$ $(4.9\pm 24.4)\cdot 10^{-24}$ $[C^{S8,RR}_{uu}]_{1221}$ $(2.7\pm 14.9)\cdot 10^{-25}$ $[C^{S8,RR}_{dd}]_{1331}$ $(1.1\pm 4.2)\cdot 10^{-24}$
$[C^{S8,RR}_{dd}]_{1111}$ $(-1.5\pm 4.6)\cdot 10^{-24}$ $[C^{V1,LR}_{uddu}]_{2112}$ $(-5.5\pm 22.3)\cdot 10^{-25}$ $[C^{S1,RR}_{dd}]_{2332}$ $(-6.4\pm 4.7)\cdot 10^{-28}$
$[C^{S1,RR}_{dd}]_{2222}$ $(-4.4\pm 2.8)\cdot 10^{-29}$ $[C^{V8,LR}_{uddu}]_{2112}$ $(-1.7\pm 10.2)\cdot 10^{-25}$ $[C^{S8,RR}_{dd}]_{2332}$ $(-1.3\pm 0.8)\cdot 10^{-27}$
$[C^{S8,RR}_{dd}]_{2222}$ $(-5.3\pm 3.2)\cdot 10^{-29}$ $[C^{V1,LR}_{uddu}]_{2222}$ $(3.6\pm 2.4)\cdot 10^{-28}$ $[C^{S1,RR}_{ud}]_{1133}$ $(7.7\pm 61.1)\cdot 10^{-25}$
$[C^{S1,RR}_{dd}]_{1122}$ $(1.3\pm 5.7)\cdot 10^{-24}$ $[C^{V8,LR}_{uddu}]_{2222}$ $(1.2\pm 0.8)\cdot 10^{-27}$ $[C^{S8,RR}_{ud}]_{1133}$ $(2.6\pm 13.2)\cdot 10^{-25}$
$[C^{S8,RR}_{dd}]_{1122}$ $(-3.7\pm 14.2)\cdot 10^{-25}$ $[C^{V1,LR}_{uu}]_{1221}$ $(-3.6\pm 21.9)\cdot 10^{-25}$ $[C^{S1,RR}_{uddu}]_{1331}$ $(-5.1\pm 24.2)\cdot 10^{-24}$
$[C^{S1,RR}_{dd}]_{1221}$ $(1.1\pm 4.1)\cdot 10^{-24}$ $[C^{V8,LR}_{uu}]_{1221}$ $(1.8\pm 6.7)\cdot 10^{-25}$ $[C^{S8,RR}_{uddu}]_{1331}$ $(5.6\pm 33.0)\cdot 10^{-25}$
$[C^{S8,RR}_{dd}]_{1221}$ $(-3.5\pm 13.1)\cdot 10^{-25}$ $[C^{S1,RR}_{ud}]_{2233}$ $(-5.1\pm 2.8)\cdot 10^{-29}$
$[C^{S1,RR}_{ud}]_{1111}$ $(-2.3\pm 3.6)\cdot 10^{-25}$ $[C^{S8,RR}_{ud}]_{2233}$ $(-1.7\pm 0.9)\cdot 10^{-29}$
$[C^{S8,RR}_{ud}]_{1111}$ $(7.0\pm 10.8)\cdot 10^{-26}$ $[C^{S1,RR}_{uddu}]_{2332}$ $(3.3\pm 2.0)\cdot 10^{-28}$
$[C^{S1,RR}_{ud}]_{1122}$ $(-1.4\pm 5.1)\cdot 10^{-24}$ $[C^{S8,RR}_{uddu}]_{2332}$ $(-4.0\pm 1.8)\cdot 10^{-29}$
$[C^{S8,RR}_{ud}]_{1122}$ $(4.5\pm 18.1)\cdot 10^{-25}$ $[C^{V1,LR}_{uddu}]_{2332}$ $(9.3\pm 4.7)\cdot 10^{-29}$
$[C^{S1,RR}_{uddu}]_{1111}$ $(-2.4\pm 3.8)\cdot 10^{-25}$ $[C^{V8,LR}_{uddu}]_{2332}$ $(-2.6\pm 1.4)\cdot 10^{-29}$
$[C^{S8,RR}_{uddu}]_{1111}$ $(7.0\pm 10.6)\cdot 10^{-26}$
$[C^{S1,RR}_{uddu}]_{1221}$ $(-1.6\pm 6.0)\cdot 10^{-24}$
$[C^{S8,RR}_{uddu}]_{1221}$ $(4.6\pm 14.6)\cdot 10^{-25}$
@@
Table 9: Master formula for $d_{\rm Hg}$ in WET in the JMS basis at the EW scale. The $\alpha_{I}^{\rm Hg}$ have units of $e$ cm $\rm TeV^{2}$ and $e$ cm $\rm TeV$ for the four-fermion and dipole operators, respectively. The $90\%$ CL upper bound on $d_{\rm Hg}$ is $6.2\times 10^{-30}$ $e$ cm.